OpenStudy (anonymous):

if you subtract 13 times a number from its square, the result is -36 translate into an equation and find the factors

OpenStudy (anonymous):

13 times a number=13x its square=x^2 so, x^2-13x=-36 do you know how to solve it from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Move the -36 over so the equation is equal to 0... x^2-13x+36=0 you've never done factoring?

OpenStudy (anonymous):

k but from there

OpenStudy (anonymous):

all i know is 9 and 4 but is one negtive positive???

OpenStudy (anonymous):

theyre both negative because -9-4=13 and -9*-4=36

OpenStudy (anonymous):

typo i meant -13

OpenStudy (anonymous):

k ty for your help

OpenStudy (anonymous):

So you got (x-9)(x-4) as your factors right?